### v1

# How Far Are We from Optimal Reasoning Efficiency?

Abstract
Large Reasoning Models (LRMs) demonstrate remarkable problem-solving capabilities through extended Chain-of-Thought (CoT) reasoning but often produce excessively verbose and redundant reasoning traces. This inefficiency incurs high inference costs and limits practical deployment. While existing fine-tuning methods aim to improve reasoning efficiency, assessing their efficiency gains remains challenging due to inconsistent evaluations. In this work, we introduce the reasoning efficiency frontiers, empirical upper bounds derived from fine-tuning base LRMs (DeepSeek-R1-Distill-Qwen-1.5B/7B and Qwen3-4B/8B) across diverse approaches and training configurations. Based on these frontiers, we propose the Reasoning Efficiency Gap (REG), a unified metric quantifying deviations of any fine-tuned LRMs from these frontiers. Systematic evaluation on challenging mathematical benchmarks reveals significant gaps in current methods: they either sacrifice accuracy for short length or still remain inefficient under tight token budgets. To reduce the efficiency gap, we propose REO-RL, a class of Reinforcement Learning algorithms that minimizes REG by targeting a sparse set of token budgets. Leveraging numerical integration over strategically selected budgets, REO-RL approximates the full efficiency objective with low error using a small set of token budgets. Through systematic benchmarking, we demonstrate that our new efficiency metric, REG, effectively captures the accuracy-length trade-off, with low-REG methods reducing length while maintaining accuracy. Our approach, REO-RL, significantly improves reasoning efficiency, reducing REG by $\geq 50\%$ across all evaluated LRMs and matching Qwen3-4B/8B efficiency frontiers under a 16K token budget with minimal accuracy loss. Ablation studies confirm the effectiveness of our exponential token budget strategy. Finally, our findings highlight that fine-tuning LRMs to perfectly align with the efficiency frontiers remains an open challenge. We will release the code, data, and models. [^1] https://github.com/samjia2000/Optimal-Reasoning-Efficiency.
1 Introduction
Large Reasoning Models (LRMs) have recently emerged as a powerful class of models capable of solving complex tasks that require advanced reasoning. Frontier LRMs such as OpenAI o1 [31] and DeepSeek R1 [9] have obtained superior performance across a wide range of tasks, including mathematical reasoning and competitive programming. A major factor behind this success is their ability to perform deep, multi-step reasoning through extended Chain-of-Thought (CoT) processes. These reasoning traces often include sophisticated operations such as reflection, verification, and exploration, within a single inference pass.
However, the powerful capability of long CoT reasoning comes at a cost. LRMs frequently generate overly verbose and redundant reasoning traces, a phenomenon referred to as the overthinking problem [51, 40]. Recent studies [4, 40] have shown that even simple questions like “2 + 3 = ?” can result in outputs spanning up to 900 tokens. This redundancy brings a significant cost in inference time and limits practical deployment. Several fine-tuning approaches have been proposed to improve reasoning efficiency, with representative examples being length reduction with RL [43, 24, 1, 2, 52, 38, 34, 50, 36, 12], and hybrid reasoning approaches that train LRMs to adaptively select thinking or no-thinking modes [15, 23, 58, 19]. However, comparing these methods remains difficult due to inconsistent experiment setups, including varying models, evaluation benchmarks, and mixed performance metrics. It is still unclear how close current approaches are to the optimal trade-off between length and accuracy.
In this work, we investigate a critical question: How far are the current approaches from reaching the optimal reasoning efficiency? To answer this question, we conduct a comprehensive empirical study using a set of LRMs, DeepSeek-Distill-Qwen-1.5B/7B and Qwen3-4B/8B, on a set of challenging mathematical reasoning benchmarks. We introduce the concept of reasoning efficiency frontiers, derived from fine-tuning the base LRMs with various types of algorithms and diverse training configurations. These reasoning efficiency frontiers represent the best reward achievable by the current approaches at each token budget, offering a practical lower bound on optimal efficiency. By comparing current methods to these frontiers, we uncover a substantial gap. Existing methods often fall short in one of two ways, either (1) they aggressively shorten responses at the expense of accuracy, or (2) methods that reach high overall accuracy would consume significantly more tokens than necessary to reach moderate accuracy levels. To quantify this gap, we propose the Reasoning Efficiency Gap (REG), a unified metric that captures both accuracy and response length by measuring the area between the budget-accuracy curve of an LRM and the efficiency frontiers. Specifically, REG is computed as the accuracy gap between the efficiency frontiers and the LRM across all token budgets. REG also quantifies the average amount of wasted tokens across all accuracies. More importantly, REG offers practical insights into how much room still remains for improvement.
We further ask: How can an LRM be fine-tuned to minimize this efficiency gap? A natural approach is to optimize the rewards across all possible token budgets at RL training time. However, this approach leads to a costly training process since rewards across all token budgets should be evaluated. To overcome the inefficiency of this dense reward approach, we introduce Reasoning Efficiency Optimization with Reinforcement Learning (REO-RL), a novel family of RL algorithms for improving reasoning efficiency of LRMs. The key insight of REO-RL is that the total rewards across all token budgets can be well-approximated using numerical integration over a small set of representative token budgets. We introduce two installations of REO-RL, including REO-RL (Oracle) that uses an oracle-based greedy approach to select token budgets based on the estimated reasoning efficiency frontiers, REO-RL (Exp) that selects exponentially spaced token budgets.
Through systematically benchmarking a diverse set of existing methods, we find that our efficiency metric REG effectively captures the trade-off between accuracy and response length. Methods that achieve low REG could significantly reduce the average response length with slight or even no accuracy drop. By reducing the efficiency gap, our approach, REO-RL, outperforms baselines in terms of reasoning efficiency, and reduces REG by at least 50% across all evaluated LRMs. Notably, REO-RL matches the efficiency frontiers of Qwen3-4B/8B under 16K token budget, with only slight overall accuracy drop. Our ablation study further validates the success of the exponential token budget selection strategy and effective approximation with a small amount of token budgets. Finally, our findings highlight an unresolved challenge: fine-tuning LRMs to precisely align with their efficiency frontiers remains an open problem.
2 Related Works
Efficient Reasoning.
Prior studies have shown that LRMs often suffer from redundant reasoning. Even for very simple questions, Frontier LRMs often generate lengthy responses spanning thousands of tokens [5, 40]. This redundancy in the reasoning process brings significant overheads in the inference costs. Several works are then proposed to make LRM reasoning more concise. [43, 24, 1, 2, 52, 38, 34, 50, 36, 12, 33] investigate RL training with length reward designs, mostly focusing on reducing the reasoning lengths. A line of works apply SFT to fine-tune LRMs on datasets with variable-length reasoning traces to elicit concise reasoning [54, 45, 10] or adjustable length control [16, 47, 27, 21, 56]. Some recent works investigate training LRMs to adaptively select whether to perform an explicit thinking process or directly produce the final solution according to the input query [58, 44, 15]. Test-time techniques are also applied for enhancing the reasoning efficiency through reward model guided decoding [41, 20], uncertainty-based dynamic reasoning [7, 8], and confidence-based approaches [42, 14]. In this work, we study the optimal reasoning efficiency for LRMs and focus on training-based approaches for enhancing LRM reasoning efficiency. Our method aims at enhancing the accuracy of the LRM under diverse token budgets without explicitly incentivizing shorter responses.
RL for LRM Reasoning.
Reinforcement Learning is the central technique for eliciting and enhancing the reasoning capability of LRMs. Frontier LRMs, including OpenAI o1 [31] and DeepSeek R1 [9], have shown that applying "zero RL" on a base LLM could effectively elicit the ability to utilize long CoTs for complex reasoning. A series of works have emerged with the focus on improving the training efficiency of RL for LRMs from the perspectives of data [26, 35, 11, 18, 46], algorithms [9, 11, 26, 55, 57], and training framework [39, 25, 35]. A number of works successfully apply zero RL training on a wide range of reasoning-heavy domains, including multi-modality [37, 59], medical [53, 3], and financial [22]. Recent works also explore efficiency enhancement by encouraging concise reasoning with RL [43, 24, 1, 2, 52, 38, 34, 50, 36, 12]. In this work, we focus on enhancing the reasoning efficiency with RL.
3 Preliminary
LRM Reasoning.
In this work, we focus on the task of mathematical reasoning. Given a question $x$, the goal of an LRM policy is to generate a response $y$ that contains step-by-step reasoning to derive the correct answer. We assume access to a verifier $\mathcal{R}(x,y)$ that evaluates the correctness of a solution $y$ given the question $x$. In practice, such a verifier is implemented by matching the ground-truth answer and the model-generated answer. The LRM is a policy $\pi_{\theta}$ parameterized with $\theta$ and generates a sequence of reasoning tokens in an auto-regressive manner. Given a question distribution $\mathcal{D}$, the objective of the LRM is to maximize the probability of producing correct responses,
$J(\mathcal{D},\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[\mathcal{R}(x,y)]$ (1)
where $\theta$ is in a parameter space $\Theta$ and the response length $|y|$ is limited to the maximum length $L_{\text{max}}$. In practice, $\theta$ is usually obtained through applying fine-tuning approaches such as RL on a base LRM. Therefore we assume the existence of a base LRM $\theta_{\text{base}}$ and $\Theta$ to be the set of all LRMs that could be obtained by fine-tuning $\theta_{\text{base}}$ with any algorithm.
4 Understanding the Limits of Efficient Reasoning
4.1 Defining Optimality in Token-Bounded Reasoning
Evaluating Reasoning under Token Budgets.
To assess optimal reasoning efficiency, we must evaluate the performance of an LRM $\pi_{\theta}$ under a fixed token budget $L$. Simply truncating an output after $L$ tokens, however, may lead to incomplete responses with no answers. To address this, we define a fallback mechanism following prior works [29, 7]. If the reasoning trace $y\sim\pi_{\theta}(\cdot|x)$ exceeds $L$ tokens, the model is prompted to produce a final answer $a$ directly from the truncated trace $y_{:L}$,
$a=\mathrm{Answer}(\pi_{\theta},x,y)=\begin{cases}\pi_{\theta}(\cdot|x,y_{:L},[\text{The Final Answer is}])\quad&\text{if}\;|y|>L\\ \mathrm{ExtractAnswer}(x,y)&\text{otherwise}\end{cases}$
where $y_{:L}$ denotes the first $L$ tokens of the reasoning trace and $\mathrm{ExtractAnswer}(x,y)$ extracts the final answer from a complete trace. This approach allows consistent evaluation across different budgets, though it introduces a minor additional token cost in the truncation case. For conciseness, we use $r(x,y_{:L};\theta)$ to represent the expected reward for $\pi_{\theta}$ on the response $y_{:L}$, i.e.
$r(x,y_{:L};\theta)=\mathcal{R}(x,\mathrm{Answer}(\pi_{\theta},x,y_{:L}))$ (2)
Length-Constrained Reward and Optimality.
We define the length-constrained reward for a model $\pi_{\theta}$ over a question distribution $\mathcal{D}$ as the expected reward obtained when the model is restricted to a token budget $L$,
$J(\mathcal{D},\theta,L)=\mathbb{E}_{x\sim\mathcal{D}}[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[r(x,y_{:L};\theta)]]$ (3)
The length-constrained optimal reward then captures the best possible reward achievable by any model in a parameter space $\Theta$ under the same budget,
$J_{\text{optimal}}(\mathcal{D},\Theta,L)=\max_{\theta\in\Theta}J(\mathcal{D},\theta,L)$ (4)
4.2 Empirical Estimation of Reasoning Frontiers
Figure 1: Reasoning Efficiency Frontiers & Reasoning Efficiency Gap. (a) Starting from a base LRM $\pi_{\theta_{\text{base}}}$, we apply diverse fine-tuning strategies to obtain a large amount of LRMs. (b) We then compute the best achievable accuracy across varying token budgets to obtain the reasoning efficiency frontiers (Eq. 5). Reasoning Efficiency Gap (REG) is a unified metric that captures both accuracy and length by measuring the area between the budget-accuracy curve of an LRM and the efficiency frontiers.
We aim to characterize the optimal reasoning efficiency, denoted by $J_{\text{optimal}}(\mathcal{D},\Theta,L)$, that reflects the best achievable reward at any token budget $L$ across all possible model parameters $\theta\in\Theta$. However, computing this optimal frontier exactly is infeasible in practice, as it requires exhaustively exploring all algorithms and training configurations. Instead, we construct an empirical reasoning efficiency frontier by fine-tuning a diverse set of models using existing approaches. Let $\hat{\Theta}=\{\theta_{1},\ldots,\theta_{m}\}\subseteq\Theta$ denote the collection of parameters from $m$ fine-tuned models. Based on these, we define the empirical reasoning efficiency frontier as,
Definition 4.1:  Reasoning Efficiency Frontier Given a parameter space $\Theta$, a set of model parameters $\hat{\Theta}=\{\theta_{1},\cdots,\theta_{m}\}$ and a question distribution $\mathcal{D}$, we define the reasoning efficiency frontier $\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$ as $\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)=\max_{\theta\in\{\theta_{1},\cdots,\theta_{m}\}}J(\mathcal{D}_{t},\theta,L)\;\;\;\;\forall L\in[1,L_{\text{max}}]$ (5)
Note that $\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$ serves as a lower bound of the optimal frontier since $\hat{\Theta}$ is a subset of $\Theta$,
$\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)\leq J_{\text{optimal}}(\mathcal{D},\Theta,L)$
Diverse Fine-tuning Approaches.
To obtain a close approximation to the optimal frontier in Eq. 5, we fine-tune models using a wide range of training strategies,
  • Online RL with Token Budgets: We conduct online RL training with different token budgets ranging from 512 to 32k. Fine-tuning an LRM with a token budget effectively enforces the LRM to reason with limited tokens [48, 12].
  • Online RL with Length Rewards. We test various reward designs that promote concise yet accurate reasoning, including length-harmonizing rewards [24] and length-group normalized rewards [2].
  • Preference Learning. We apply SimPO [28] with preference datasets constructed via methods such as TOPS [51] and DAST [38]. For example, we contrast short correct responses with longer ones to promote conciseness [30].
Experiment Setup.
We conduct our study using two classes of base LRMs, DeepSeek-R1-Distill-Qwen-1.5B/7B [9] and Qwen3-4B/8B [49]. Specially, for DeepSeek-R1-Distill-Qwen-1.5B/7B, we start from RL-fine-tuned versions and further fine-tune them using the strategies above. The evaluation is carried out on a set of challenging mathematical reasoning benchmarks. More training and evaluation details can be found in Sec. 6.1 and Appendix. C.
Reasoning Efficiency Frontiers.
Based on the fine-tuned models, we are able to construct the reasoning efficiency frontiers for the base LRMs, as illustrated in Fig. 2. Interestingly, we find that DeepSeek-R1-Distill-Qwen models are both far from the efficiency frontiers, with clear efficiency gaps. By contrast, Qwen3 models reach a similar accuracy as the efficiency frontiers and are much closer to the efficiency frontiers, allowing limited improvement space for reasoning efficiency.
(a) DeepSeek-R1-Distill-Qwen-1.5B
(b) DeepSeek-R1-Distill-Qwen-7B
(c) Qwen3-4B
(d) Qwen3-8B
Figure 2: Reasoning efficiency frontiers for DeepSeek-R1-Distill-Qwen-1.5B/7B and Qwen3-4B/8B. Comparing the two classes of models, Qwen3 models have higher accuracy and are closer to the efficiency frontiers than DeepSeek-R1-Distill-Qwen models.
Measuring the Gap to Optimality.
Given the estimated reasoning efficiency frontiers $\hat{J}_{\text{optimal}}(\mathcal{D},\Theta,L)$, we introduce the metric of Reasoning Efficiency Gap (REG) to quantify the distance of any LRM from reaching the optimal reasoning efficiency.
Definition 4.2:  Reasoning Efficiency Gap (REG) Given any LRM $\pi_{\theta}$ and the estimated reasoning efficiency frontier $\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$, we define Reasoning Efficiency Gap as, $d_{\text{REG}}(\theta,\mathcal{D},\hat{\Theta})=\sum_{L=1}^{L_{\text{max}}}\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)-J(\mathcal{D},\theta,L)$ (6)
Note that, when $\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$ and $J(\mathcal{D},\theta,L)$ are both non-decreasing, REG could equivalently represent the average amount of wasted tokens across all accuracies.
Advantages of REG.
As a single metric, REG unifies both the accuracy and length aspects. At the same time, REG quantitatively assesses the gap of reasoning efficiency between an LRM and efficiency frontiers. Most of the existing evaluation metrics only focus on one of the accuracy and length aspects. For example, Pu et al. [32] proposes to evaluate reasoning redundancy by computing the length difference between the longest and the shortest response, without considering correctness of the generated responses. We also note that Chen et al. [4] introduces an outcome efficiency metric that measures the ratio between the position of first reaching the correct answer and the the overall length for correct responses. Although this outcome efficiency metric takes both accuracy and length into consideration, it has the risk of being hacked if the LRM only produces an answer at the end of the response.
5 Methodology
5.1 Boosting Reasoning Efficiency by Optimizing Length-Constrained Rewards
Optimizing Length-Constrained Rewards.
To minimize the reasoning efficiency gap, a straightforward idea is to optimize the length-constrained rewards under all token budgets to enhance the reasoning efficiency of $\pi_{\theta_{\tt{base}}}$, leading to the efficiency objective,
$\mathscr{L}_{\text{Efficiency}}(\theta,\mathcal{D})=\sum_{L=1}^{L_{\tt{max}}}J(\mathcal{D},\theta,L)$ (7)
where $L_{\tt{max}}$ is the maximum generation length. However, directly optimizing Eq. 7 is computationally impractical. Evaluating $J(\mathcal{D},\theta,L)$ for each budget $L\in[1,L_{\tt{max}}]$ requires separate inference runs to evaluate truncated responses as discussed in Sec. 4.1. As a result, each training example would require up to $L_{\tt{max}}$ additional LRM generations, leading to increased compute time and memory usage.
5.2 Reasoning Efficiency Optimization with Reinforcement Learning
REO-RL.
We introduce Reasoning Efficiency Optimization with Reinforcement Learning (REO-RL), a class of an efficient training algorithms that optimize an approximation for the objective in Eq. 7. In REO-RL, instead of optimizing the length-constrained reward in all token budgets, we approximate the objective with a small set of selected token budgets $L_{1},\cdots,L_{N}$ to ensure high training efficiency. Specifically, following the Trapezoidal rule in numerical integration, we could approximate the objective in Eq. 7 with,
$\sum_{L=1}^{L_{\tt{max}}}J(\mathcal{D},\theta,L)$ $\approx\sum_{i=1}^{N}\frac{L_{i+1}-L_{i-1}}{2}J(\mathcal{D},\theta,L_{i})+\frac{L_{1}}{2}\cdot J(\mathcal{D},\theta,0)+\frac{L_{\text{max}}-L_{N}}{2}\cdot J(\mathcal{D},\theta,L_{\text{max}})$ (8)
$=f(\mathcal{D},\theta,\{L_{1},\cdots,L_{N}\})$ (9)
where we assume $L_{0}=0$ and $L_{N+1}=L_{\text{max}}$ and $f(\mathcal{D},\theta,\{L_{1},\cdots,L_{N}\})$ denotes the approximated objective with token budgets $L_{1},\cdots,L_{N}$. As the number of selected token budgets $N$ increases, $f(\mathcal{D},\theta,\{L_{1},\cdots,L_{N}\})$ would become closer and closer to the objective in Eq. 7.
The approximated objective $f(\mathcal{D},\theta,\{L_{1},\cdots,L_{N}\})$ could be equivalently represented as an RL objective with dense rewards, leading to the objective of REO-RL,
REO-RL: $\mathscr{L}_{{REO-RL}}(\theta,\mathcal{D})=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[\sum_{i=1}^{N+1}c_{i}\cdot r(x,y_{:L_{i}};\theta)\right]\right]$ (10)
where $c_{i}=\frac{L_{i+1}-L_{i-1}}{2}$ for $1\leq i\leq N$ and $c_{N+1}=\frac{L_{\text{max}}-L_{N}}{2}$ are the coefficient for the $i$-th token budget following Eq. 9.
Figure 3: Selection of Token Budgets in REO-RL. (a) Token budgets selected roughly follow an exponential pattern. (b) Both the oracle greedy approach and the exponential approach achieve lower approximation errors with a few token budgets.
Different sets of token budgets $L_{1},\cdots,L_{N}$ would cause different approximation error in Eq. 9. Ideally, the approximation error induced by the selected token budgets should be low to ensure that REO-RL aligns with the original objective in Eq. 5.
REO-RL (Oracle).
Note that the original objective in Eq. 5 is bounded by the theoretical reasoning efficiency frontier in Sec. 4.2, i.e. $\sum_{L=1}^{L_{\text{max}}}J(\mathcal{D},\theta,L)\leq\sum_{L=1}^{L_{\text{max}}}J_{\text{optimal}}(\mathcal{D},\Theta,L)$. A natural idea is to determine the optimal token budget selection scheme based on the estimated reasoning efficiency frontiers. Specifically, for any set of token budgets $L_{1},L_{2},\cdots,L_{N}$, $\sum_{L=1}^{L_{\text{max}}}\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$ could be approximated in a similar way as Eq. 9,
$f_{\text{optimal}}(\mathcal{D},\hat{\Theta},\{L_{1},\cdots,L_{N}\})$ $=\sum_{i=1}^{N}\frac{L_{i+1}-L_{i-1}}{2}\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L_{i})+\frac{L_{1}}{2}\cdot\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},0)$ (11)
$+\frac{L_{\text{max}}-L_{N}}{2}\cdot\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L_{\text{max}})$ (12)
where $f_{\text{optimal}}(\mathcal{D},\hat{\Theta},\{L_{1},\cdots,L_{N}\})$ denotes the approximated value for $\sum_{L=1}^{L_{\text{max}}}\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)$ given token budgets $L_{1},\cdots,L_{N}$.
We adopt a greedy approach that iteratively selects the token budget with the lowest approximation error. Note that this is an oracle approach since the reasoning efficiency frontiers should be known in advance. This greedy selection approach leads to an oracle algorithm, REO-RL (Oracle). In REO-RL (Oracle), the token budgets $L_{1},\cdots,L_{N}$ are selected according to,
$L_{i}=\arg\min_{L^{\prime}}\mid f_{\text{optimal}}(\mathcal{D},\hat{\Theta},\{L_{1},\cdots,L_{i-1},L^{\prime}\})-\sum_{L=1}^{L_{\text{max}}}\hat{J}_{\text{optimal}}(\mathcal{D},\hat{\Theta},L)\mid$
The oracle greedy approach could produce a set of token budgets that achieves low approximation error in Fig. 3. As illustrated in Fig. 3, the approximation error gradually degrades with more token budgets. Notably, the approximation error could be lower than $1\%$ with $N\geq 5$.
REO-RL (Exp).
However, in cases when the reasoning efficiency frontiers are unknown, it is infeasible to apply REO-RL (Oracle). We observe that token budgets selected by the greedy selection approach roughly follow an exponentially spaced pattern as shown in Fig. 3. Therefore, we propose to adopt an exponentially spaced scheme for token budget selection, leading to the algorithm, REO-RL (Exp), that selects a set of exponentially spaced token budgets,
$L_{i}=L_{\text{min}}\cdot\left({L_{\text{max}}}/{L_{\text{min}}}\right)^{\frac{i-1}{N}}$
where $L_{\text{min}}/L_{\text{max}}$ are the minimum/maximum token budgets.
6 Experiments
6.1 Experimental Setup
Models, Datasets & Metrics.
We use DeepSeek-R1-Distill-Qwen-1.5B&7B and Qwen3-4B&8B as the base LRMs. For training, we adopt a mixture of training data consisting of 135k problems sourced from DeepScaleR [26] and AReaL [35] For evaluation, we use several challenging mathematical benchmarks: AMC 2023, AIME 2024 & 2025, and Minerva Math [17]. We report the average accuracy of 32 responses generated with temperature $T=0.6$ and top_p$=0.95$ with maximum length $L_{\max}=32K$. The main results are averaged over three benchmarks. When evaluating REG, we use a smaller length $L_{\text{max}}=16K$.
REO-RL & Baselines.
For REO-RL, besides REO-RL (Exp) and REO-RL (Oracle), we further consider REO-RL (Q-Spec), a variant of REO-RL that optimizes rewards in question-specific minimum reasoning token budgets [^2] For more details on REO-RL (Q-Spec), please refer to Appendix E.
  • Online RL with Length-based Rewards: We consider online RL with length-based rewards, including length-harmonizing rewards [24], and length group normalized rewards [2]. We make comparison with Meta Reinforcement Fine-tuning (MRT) [34], which minimizes regret for individual steps to enhance reasoning efficiency. Additionally, we also adopt online RL with hard token budgets as a baseline, which is also adopted in recent works [12, 48]. We note that a recent work BRPO [33] optimizes rewards under linearly spaced token budgets, which is similar to a variant of REO-RL studied in the ablation study (Sec. 6.4). [^3] A more detailed discussion between BRPO and REO-RL is provided in Appendix. A
  • Hybrid Reasoning. We use HGPO [15] as a representative hybrid reasoning baseline that trains the LRM to adaptively select between the thinking and no-thinking modes. There are also recent works sharing the similar idea of encouraging no-thinking mode when no-thinking mode achieves competitive accuracy [6, 58, 44].
  • Supervised Fine-Tuning: For each problem in the training data, we generate multiple responses and perform SFT on the shortest correct one. Two strategies are used for data generation. The first strategy is direct generation with the problems as inputs [30]. The second strategy follows TOPS to prompt the LRM with different levels of reasoning efforts [51].
  • Preference Learning: We apply SimPO [28] on various preference datasets. In $SimPO_{shortest}$, we use the shortest correct response and the longest response as the preference pairs. We also follow TOPS [51] and DAST [38] to construct preference datasets.
We have also tried RL-based length control methods such as  [1, 48] but find these methods only achieve successful length control under low token budgets and yield similar length-accuracy trade-off as RL w. Token Budgets. More details about the baselines and our other investigations could be found in Appendix. D.
(a) DeepSeek-R1-Distill-Qwen-1.5B
(b) DeepSeek-R1-Distill-Qwen-7B
(c) Qwen3-4B
(d) Qwen3-8B
Figure 4: Performance comparison of REO-RL (Exp) with representative baseline methods under 16K token budget. REO-RL (Exp) outperforms baselines in terms of reasoning efficiency. For DeepSeek-R1-Distill-Qwen models, there still exist large performance gaps under tight token constraints. For Qwen3 models, though REO-RL matches the efficiency frontiers under the 16K token budget, accuracy loss still exists (See Tab. 1). Results are averaged over AMC 2023, AIME 2024 & 2025, and Minerva Math.
Figure 5: Reasoning efficiency frontiers of DeepSeek-R1-Distill-Qwen-1.5B/7B and Qwen3-4B/8B. Qwen3 models obtain higher accuracy. Interestingly, DeepSeek-R1-Distill-Qwen-7B could reason more efficiently than Qwen3 models under the 4K token budget.
Training Details.
For REO-RL and all baselines, we implement all methods based on the AReaL framework [35]. For DeepSeek-R1-Distill-Qwen-1.5B/7B, we fine-tune the corresponding RL-trained versions. Specifically, we use AReaL-Boba-RL-1.5B [35] and SkyWork-OR1-Math-7B [11] as the starting points for further fine-tuning in 1.5B and 7B experiments, respectively. For Qwen3 models, we directly apply RL training on the base LRMs. For REO-RL (Exp) and REO-RL (Oracle), we use $N=5$ token budgets since $N=5$ already obtains sufficiently accurate approximations as illustrated in Fig. 3. For more training details, please refer to Appendix. C.
6.2 Main Results
Figure 6: REG effectively captures the trade-off between accuracy and response length. Achieving a low REG require both competitive accuracy and short response length. By minimizing the efficiency gap, REO-RL outperforms baselines in terms of reasoning efficiency.
Fig. 4 presents the budget-accuracy curves of LRMs fine-tuned using REO-RL (Exp) and a range of representative baseline methods. [^4] For clearer visualization, we include only a set of representative approaches and plot their performance under limited token budgets. Fig. 6 further illustrates the accuracy, response length, and REG reduction for different classes of methods. Full quantitative results are detailed in Tab. 1. We highlight several key conclusions,
There exist fundamental gaps between existing approaches and the frontiers. As shown in Fig. 4, several baselines can approach or even reach the reasoning efficiency frontier given sufficient token budgets. However, they could require significantly more tokens than the efficiency frontier to achieve moderate-level accuracies. On the other hand, some methods, such as RL with Token Budget, perform well under tight token constraints but exhibit substantially lower overall accuracy. Notably, our proposed approach, REO-RL, could match the efficiency frontiers under low token budget for Qwen3 models, but still with slight drop in the overall accuracy, as illustrated in Tab. 1. Our benchmarking results suggest that optimizing the base LRM to precisely match reasoning efficiency frontiers remains an open problem.
REG effectively captures the trade-off between accuracy and response length. As shown in Fig. 6, achieving low REG requires balancing high accuracy with concise output. For example, REO-RL maintains strong accuracy while generating compact responses, resulting in significant reduction in REG. In contrast, methods that optimize only one aspect, such as minimizing length at the cost of accuracy, fail to achieve low REG. RL with Token Budgets, despite producing shorter responses, suffers from significant accuracy loss.
Finally, REO-RL consistently outperforms baselines in terms of reasoning efficiency. Compared to baselines, the of REO-RL are much closer to the reasoning efficiency frontiers, exhibiting much more reduction in the REG. REO-RL maintains high accuracy while generating shorter outputs. REO-RL (Exp) consistently reduces the REG by at least 50% across all base LRMs. For Qwen3 models, baselines that apply RL with length-based rewards could not consistently reduce the efficiency gap. In fact, we find that RL w. Length Harmonizing Rewards even enlarges the efficiency gap for Qwen3-4B/8B models.
Method AIME 2024 AIME 2025 AMC 2023 Minerva Math Average
Acc (%) $\uparrow$ Len $\downarrow$ REG $\downarrow$ Acc (%) $\uparrow$ Len $\downarrow$ REG $\downarrow$ Acc (%) $\uparrow$ Len $\downarrow$ REG $\downarrow$ Acc (%) $\uparrow$ Len $\downarrow$ REG $\downarrow$ Acc (%) $\uparrow$ Len $\downarrow$ REG $\downarrow$
DeepSeek-R1-Distill-Qwen-1.5B
Base LRM 29.2 16757.1 2239.0 23.5 16577.6 1485.9 71.6 9956.8 2425.1 32.0 8212.0 970.7 39.1 12875.9 1780.2
Vanilla RL 42.2 12902.2 1266.2 28.2 13984.8 1077.7 81.9 8044.4 1742.6 35.2 8305.8 641.3 46.9 10809.3 1181.9
REO-RL (Exp) (ours) 42.9↑0.7 8780.3 $\downarrow{46.6\%}$ 31.9↑3.6 8630.5 $\downarrow{64.4\%}$ 84.5↑2.7 4990.8 $\downarrow{63.6\%}$ 36.1↑0.8 5094.7 $\downarrow{51.7\%}$ 48.8↑2.0 6874.1 $\downarrow{57.6\%}$
REO-RL (Oracle) (ours) 42.5↑0.3 8443.8 $\downarrow{30.0\%}$ 34.7↑6.5 8288.7 $\downarrow{85.4\%}$ 84.4↑2.5 4626.6 $\downarrow{53.1\%}$ 35.8↑0.5 4427.8 $\downarrow{56.5\%}$ 49.3↑2.5 6446.7 $\downarrow{54.7\%}$
REO-RL (Q-Spec) (ours) 46.6↑4.4 9162.3 $\downarrow{90.7\%}$ 32.7↑4.5 8832.4 $\downarrow{81.4\%}$ 84.1↑2.2 5338.5 $\downarrow{66.6\%}$ 35.4↑0.2 5979.7 $\downarrow{45.0\%}$ 49.7↑2.8 7328.2 $\downarrow{73.5\%}$
RL w. Token Budget=1K 17.3↓24.9 1487.4 $\uparrow{170.4\%}$ 13.5↓14.7 1282.8 $\uparrow{140.7\%}$ 65.6↓16.2 $1$.3 $\uparrow{43.1\%}$ 32.0↓3.2 1108.8 $\downarrow{24.1\%}$ 32.1↓14.8 1247.6 $\uparrow{90.3\%}$
RL w. Token Budget=2K 22.4↓19.8 2423.9 $\uparrow{114.2\%}$ 17.5↓10.7 1966.0 $\uparrow{90.6\%}$ 73.1↓8.8 1547.1 $\downarrow{13.6\%}$ 33.9↓1.3 1671.8 $\downarrow{57.0\%}$ 36.7↓10.1 1902.2 $\uparrow{38.5\%}$
RL w. Token Budget=4K 31.4↓10.8 3991.0 $\uparrow{25.8\%}$ 26.1↓2.1 3770.8 $\downarrow{17.3\%}$ 80.9↓0.9 2561.2 $\downarrow{60.3\%}$ 35.20.0 2942.8 $\downarrow{65.5\%}$ 43.4↓3.5 3316.5 $\downarrow{28.2\%}$
RL w. Len Group Norm. Rew. 44.4↑2.2 10486.0 $\downarrow{37.5\%}$ 30.7↑2.5 10599.6 $\downarrow{33.7\%}$ 82.7↑0.9 5911.5 $\downarrow{30.7\%}$ 35.8↑0.6 6057.2 $\downarrow{34.6\%}$ 48.4↑1.5 8263.6 $\downarrow{33.3\%}$
RL w. Len-Harmonizing Rew. 41.5↓0.7 9119.2 $\downarrow{31.3\%}$ 31.6↑3.3 8113.9 $\downarrow{60.0\%}$ 85.1↑3.2 3993.2 $\downarrow{65.1\%}$ 35.4↑0.2 3721.0 $\downarrow{69.6\%}$ 48.4↑1.5 6236.8 $\downarrow{55.5\%}$
MRT 42.9↑0.7 9797.8 $\downarrow{45.3\%}$ 29.8↑1.6 9823.0 $\downarrow{41.1\%}$ 84.9↑3.0 5607.8 $\downarrow{54.2\%}$ 35.0↓0.2 5445.0 $\downarrow{26.1\%}$ 48.2↑1.3 7668.4 $\downarrow{45.0\%}$
$\text{SFT}_{\text{Shortest}}$ 43.2↑1.0 12898.1 $\downarrow{11.1\%}$ 29.9↑1.7 14221.8 $\downarrow{22.3\%}$ 82.5↑0.6 7513.7 $\downarrow{9.6\%}$ 35.2↓0.1 7641.9 $\downarrow{11.5\%}$ 47.7↑0.8 10568.9 $\downarrow{13.2\%}$
$\text{SFT}_{\text{TOPS}}$ 41.8↓0.4 10823.9 $\downarrow{8.4\%}$ 30.7↑2.5 11720.5 $\downarrow{33.7\%}$ 82.7↑0.9 6454.1 $\downarrow{17.5\%}$ 31.8↓3.4 4666.3 $\uparrow{30.4\%}$ 46.8↓0.1 8416.2 $\downarrow{12.3\%}$
$\text{SimPO}_{\text{DAST}}$ 22.9↓19.3 14029.4 $\uparrow{112.1\%}$ 19.4↓8.9 11161.7 $\uparrow{70.0\%}$ 68.1↓13.8 5955.5 $\uparrow{42.2\%}$ 32.6↓2.6 5906.4 $\uparrow{2.1\%}$ 35.7↓11.1 9263.3 $\uparrow{61.8\%}$
$\text{SimPO}_{\text{Shortest}}$ 35.8↓6.4 7422.1 $\uparrow{0.5\%}$ 26.5↓1.8 7380.9 $\uparrow{0.1\%}$ 77.7↓4.2 4020.8 $\downarrow{17.3\%}$ 34.8↓0.4 3437.0 $\downarrow{51.5\%}$ 43.7↓3.2 5565.2 $\downarrow{13.2\%}$
$\text{SimPO}_{\text{TOPS}}$ 15.6↓26.6 2647.4 $\uparrow{195.5\%}$ 18.3↓9.9 2412.2 $\uparrow{81.3\%}$ 58.5↓23.4 1497.8 $\uparrow{115.5\%}$ 25.7↓9.5 1022.8 $\uparrow{129.4\%}$ 29.5↓17.3 1895.1 $\uparrow{131.0\%}$
DeepSeek-R1-Distill-Qwen-7B
Base LRM 55.3 13062.1 1887.4 39.7 14241.9 1527.0 90.9 6177.3 1611.4 43.1 5575.8 855.7 57.2 9764.3 1470.4
Vanilla RL 66.2 14264.7 1579.0 52.9 16305.3 1453.1 93.9 7259.8 1722.2 44.6 7300.7 945.2 64.4 11282.6 1424.9
REO-RL (Exp) (ours) 64.0↓2.3 7671.5 $\downarrow{60.7\%}$ 48.8↓4.2 8361.1 $\downarrow{57.3\%}$ 93.4↓0.5 4144.6 $\downarrow{48.4\%}$ 45.5↑0.9 3687.0 $\downarrow{59.4\%}$ 62.9↓1.5 5966.0 $\downarrow{55.9\%}$
REO-RL (Oracle) (ours) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 $4$.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
REO-RL (Q-Spec) (ours) 63.9↓2.4 8407.8 $\downarrow{62.0\%}$ 51.4↓1.6 9298.5 $\downarrow{75.3\%}$ 93.4↓0.5 4230.6 $\downarrow{54.5\%}$ 43.9↓0.7 3581.0 $\downarrow{38.9\%}$ 63.1↓1.3 6379.4 $\downarrow{59.3\%}$
RL w. Token Budget=1K 26.7↓39.6 1242.4 $\uparrow{176.4\%}$ 19.7↓33.2 1168.0 $\uparrow{137.4\%}$ 73.0↓20.9 960.4 $\uparrow{59.5\%}$ 40.6↓3.9 690.7 $\downarrow{31.1\%}$ 40.0↓24.4 1015.4 $\uparrow{98.0\%}$
RL w. Token Budget=2K 36.5↓29.8 1929.6 $\uparrow{91.2\%}$ 27.6↓25.3 1831.6 $\uparrow{61.0\%}$ 84.3↓9.6 1427.5 $\downarrow{36.6\%}$ 43.9↓0.7 1210.1 $\downarrow{74.8\%}$ 48.1↓16.3 1599.7 $\uparrow{17.3\%}$
RL w. Token Budget=4K 48.3↓17.9 3480.9 $\uparrow{0.8\%}$ 35.4↓17.5 3345.2 $\downarrow{2.0\%}$ 89.1↓4.8 2109.2 $\downarrow{59.7\%}$ 43.5↓1.1 1772.8 $\downarrow{55.3\%}$ 54.1↓10.3 2677.0 $\downarrow{27.5\%}$
RL w. Len Group Norm. Rew. 64.2↓2.1 10608.1 $\downarrow{28.8\%}$ 50.8↓2.1 11486.8 $\downarrow{47.9\%}$ 94.5↑0.5 4788.9 $\downarrow{43.7\%}$ 45.2↑0.6 4309.4 $\downarrow{45.5\%}$ 63.7↓0.7 7798.3 $\downarrow{40.9\%}$
RL w. Len-Harmonizing Rew. 65.5↓0.7 9167.7 $\downarrow{37.0\%}$ 51.1↓1.8 10394.6 $\downarrow{46.1\%}$ 94.5↑0.5 4307.9 $\downarrow{52.7\%}$ 44.4↓0.2 3617.9 $\downarrow{42.2\%}$ 63.9↓0.5 6872.0 $\downarrow{44.9\%}$
MRT 66.1↓0.1 9210.6 $\downarrow{42.4\%}$ 50.5↓2.4 10559.3 $\downarrow{55.9\%}$ 93.8↓0.2 4986.6 $\downarrow{38.6\%}$ 44.7↑0.1 4852.7 $\downarrow{32.4\%}$ 63.8↓0.6 7402.3 $\downarrow{43.0\%}$
$\text{SFT}_{\text{Shortest}}$ 67.7↑1.5 13197.2 $\downarrow{13.7\%}$ 53.6↑0.7 15208.6 $\downarrow{11.1\%}$ 93.6↓0.3 6728.2 $\downarrow{4.0\%}$ 44.6↓0.0 6623.5 $\downarrow{9.4\%}$ 64.9↑0.5 10439.4 $\downarrow{9.4\%}$
$\text{SFT}_{\text{TOPS}}$ 66.0↓0.2 13204.1 $\downarrow{9.0\%}$ 52.0↓0.9 15170.8 $\downarrow{11.7\%}$ 93.3↓0.6 6272.5 $\downarrow{11.9\%}$ 44.2↓0.4 5759.7 $\downarrow{30.4\%}$ 63.9↓0.6 10101.8 $\downarrow{14.1\%}$
$\text{SimPO}_{\text{DAST}}$ 60.1↓6.1 7500.5 $\downarrow{39.0\%}$ 46.2↓6.7 7916.9 $\downarrow{47.9\%}$ 92.0↓1.9 4051.7 $\downarrow{46.4\%}$ 45.5↑0.9 3402.6 $\downarrow{61.8\%}$ 61.0↓3.4 5717.9 $\downarrow{47.3\%}$
$\text{SimPO}_{\text{Shortest}}$ 65.5↓0.7 9348.1 $\downarrow{38.5\%}$ 50.8↓2.1 10292.3 $\downarrow{45.7\%}$ 93.4↓0.5 4793.1 $\downarrow{36.2\%}$ 45.4↑0.8 4276.4 $\downarrow{46.7\%}$ 63.8↓0.6 7177.5 $\downarrow{41.0\%}$
HGPO 68.1 ↑1.9 13556.8 $\downarrow{4.6\%}$ 53.1↑0.2 15609.8 $\uparrow{2.3\%}$ 93.4↓0.5 7077.3 $\uparrow{2.7\%}$ 45.0↑0.4 6873.4 $\downarrow{15.6\%}$ 64.9↑0.5 10779.3 $\downarrow{0.4\%}$
Qwen3-4B
Base LRM 72.1 14808.7 614.3 64.7 17618.7 335.0 97.0 8397.1 955.8 46.5 6864.3 654.3 70.1 11922.2 639.9
REO-RL (Exp) (ours) 70.0↓2.1 10799.0 $\downarrow{87.8\%}$ 59.9↓4.8 13063.4 $\downarrow{45.6\%}$ 93.8↓3.2 5307.7 $\downarrow{74.4\%}$ 45.2↓1.3 3653.3 $\downarrow{40.7\%}$ 67.2↓2.8 8205.9 $\downarrow{65.2\%}$
REO-RL (Q-Spec) (ours) 67.9↓4.2 12066.1 $\downarrow{21.7\%}$ 62.1↓2.6 14755.0 $\downarrow{48.5\%}$ 95.1↓2.0 5662.4 $\downarrow{63.6\%}$ 46.0↓0.5 4203.5 $\downarrow{39.7\%}$ 67.8↓2.3 9171.7 $\downarrow{45.5\%}$
RL w. Token Budget=1K 25.9↓46.1 22992.1 $\uparrow{576.6\%}$ 23.5↓41.1 21087.8 $\uparrow{815.4\%}$ 73.6↓23.4 19078.4 $\uparrow{139.6\%}$ 43.5↓3.0 10106.0 $\uparrow{27.9\%}$ 41.6↓28.4 18316.1 $\uparrow{304.4\%}$
RL w. Token Budget=2K 39.5↓32.6 12996.9 $\uparrow{311.4\%}$ 30.1↓34.6 12546.4 $\uparrow{575.7\%}$ 85.6↓11.4 7827.0 $\downarrow{0.8\%}$ 46.0↓0.6 5542.0 $\downarrow{34.5\%}$ 50.3↓19.8 9728.1 $\uparrow{141.0\%}$
RL w. Token Budget=4K 52.6↓19.5 7909.6 $\uparrow{77.3\%}$ 37.3↓27.4 8809.0 $\uparrow{370.7\%}$ 90.8↓6.2 5316.3 $\downarrow{62.6\%}$ 47.3↑0.8 4302.1 $\downarrow{66.5\%}$ 57.0↓13.1 6584.2 $\uparrow{26.7\%}$
RL w. Token Budget=8K 64.1↓8.0 10240.9 $\downarrow{3.7\%}$ 54.2↓10.5 11042.7 $\uparrow{64.7\%}$ 93.0↓4.0 6050.0 $\downarrow{23.2\%}$ 47.4↑0.8 5067.9 $\downarrow{40.7\%}$ 64.7↓5.4 8100.4 $\downarrow{11.5\%}$
RL w. Len Group Norm. Rew. 70.0↓2.1 15022.9 $\uparrow{51.0\%}$ 63.1↓1.6 16825.9 $\uparrow{104.7\%}$ 95.6↓1.4 7682.7 $\downarrow{7.5\%}$ 47.5↑1.0 6322.9 $\downarrow{23.7\%}$ 69.1↓1.0 11463.6 $\uparrow{17.1\%}$
RL w. Len-Harmonizing Rew. 67.7↓4.4 15265.5 $\downarrow{0.4\%}$ 56.2↓8.4 18466.5 $\uparrow{183.1\%}$ 95.0↓2.0 7380.8 $\downarrow{41.9\%}$ 45.5↓1.1 7771.8 $\uparrow{15.9\%}$ 66.1↓4.0 12221.1 $\uparrow{12.3\%}$
HGPO 72.4↑0.3 15123.8 $\downarrow{16.4\%}$ 65.4↑0.7 17951.8 $\uparrow{49.6\%}$ 96.2↓0.9 8858.2 $\uparrow{16.1\%}$ 47.7↑1.2 6797.4 $\downarrow{11.7\%}$ 70.4↑0.3 12182.8 $\uparrow{5.6\%}$
Qwen3-8B
Base LRM 75.1 15221.2 1402.2 67.2 17682.8 656.1 94.8 9050.7 1026.9 48.3 7281.1 682.0 71.3 12308.9 941.8
REO-RL (Exp) (ours) 73.6↓1.5 9285.4 $\downarrow{69.7\%}$ 59.0↓8.2 11343.1 $\downarrow{68.2\%}$ 95.1↑0.3 4699.7 $\downarrow{98.6\%}$ 47.8↓0.5 3671.8 $\downarrow{62.1\%}$ 68.9↓2.5 7250.0 $\downarrow{75.9\%}$
REO-RL (Q-Spec) (ours) 73.2↓1.9 10852.9 $\downarrow{37.1\%}$ 64.4↓2.8 12645.2 $\downarrow{84.8\%}$ 94.6↓0.2 5416.1 $\downarrow{71.5\%}$ 46.8↓1.6 3952.7 $\downarrow{36.1\%}$ 69.7↓1.6 8216.7 $\downarrow{54.6\%}$
RL w. Token Budget=1K 32.3↓42.8 5170.2 $\uparrow{201.8\%}$ 22.7↓44.5 5291.7 $\uparrow{439.9\%}$ 67.6↓27.2 2042.4 $\uparrow{212.6\%}$ 39.6↓8.8 1419.3 $\uparrow{77.6\%}$ 40.5↓30.8 3480.9 $\uparrow{223.7\%}$
RL w. Token Budget=2K 52.2↓22.9 5151.0 $\uparrow{33.1\%}$ 35.1↓32.1 4696.1 $\uparrow{203.0\%}$ 84.8↓9.9 3840.4 $\uparrow{8.8\%}$ 47.4↓0.9 3034.1 $\downarrow{56.2\%}$ 54.9↓16.5 4180.4 $\uparrow{39.9\%}$
RL w. Token Budget=4K 62.3↓12.8 6905.4 $\downarrow{18.2\%}$ 45.3↓21.9 6920.0 $\uparrow{49.2\%}$ 89.0↓5.8 4274.1 $\downarrow{32.6\%}$ 48.1↓0.2 3426.4 $\downarrow{62.0\%}$ 61.2↓10.2 5381.5 $\downarrow{18.3\%}$
RL w. Token Budget=8K 71.7↓3.4 7363.9 $\downarrow{84.8\%}$ 51.4↓15.8 8359.4 $\uparrow{9.5\%}$ 92.1↓2.7 4437.3 $\downarrow{73.4\%}$ 48.4↑0.0 3534.8 $\downarrow{74.2\%}$ 65.9↓5.5 5923.8 $\downarrow{63.4\%}$
RL w. Len Group Norm. Rew. 74.2↓0.9 12841.8 $\downarrow{20.2\%}$ 64.8↓2.4 14991.5 $\downarrow{63.3\%}$ 94.5↓0.2 6366.9 $\downarrow{48.0\%}$ 46.7↓1.6 5138.1 $\downarrow{10.9\%}$ 70.1↓1.3 9834.6 $\downarrow{33.6\%}$
RL w. Len-Harmonizing Rew. 74.8↓0.3 15489.7 $\uparrow{17.7\%}$ 65.8↓1.4 18779.6 $\uparrow{54.8\%}$ 95.3↑0.5 8655.2 $\downarrow{1.1\%}$ 47.9↓0.5 7707.5 $\uparrow{17.2\%}$ 71.0↓0.4 12658.0 $\uparrow{18.9\%}$
HGPO 75.3↑0.2 15547.2 $\uparrow{10.6\%}$ 68.0↑0.8 18112.5 $\uparrow{7.6\%}$ 95.3↑0.5 9292.2 $\uparrow{9.7\%}$ 48.9↑0.5 7605.6 $\downarrow{5.9\%}$ 71.9↑0.5 12639.4 $\uparrow{6.8\%}$
Table 1: Accuracy, Generation Length, and Reasoning Efficiency Gap (REG) for all methods. For accuracy and REG, the relative changes compared to the vanilla RL baseline or the base LRMs are reported. REO-RL could significantly reduce the gap from the LRM to the reasoning efficiency frontier with slight or even no accuracy drop at the same time.
Besides these major findings, we also have the following observations,
Vanilla RL does not yield consistent improvements across budgets. Although Vanilla RL improves overall accuracy relative to the base LRM, it fails to provide stable gains across different token budgets. In the 7B experiments, its accuracy is lower than the base LRM when the token budget is below 8K.
HGPO does not necessarily improve reasoning efficiency with adaptive thinking mode switch. We find HGPO could enhance the overall accuracy but not necessarily improves the reasoning efficiency and reduces response length. In our experiments, we observe that the LRM would learn to achieve higher rewards in the no-thinking mode, along with an increasing response length during the training process of HGPO. This indicates the LRM reverts back to the thinking mode even with a no-thinking prompt, i.e. performing long-CoT reasoning outside of the “<think>…</think>” tags.
6.3 Comparison with Frontier LRMs
Claude Sonnet 3.7 (Thinking) DeepSeek R1 Qwen3-4B Qwen3-8B Qwen3-32B Vanilla RL - 7B REO-RL (Exp) - 7B (ours)
Length 17478.87 5156.39 8631.61 8898.57 7755.99 7732.11 4524.02
Accuracy 90.0% 98.2% 95.8% 94.9% 97.2% 93.5% 93.5%
Table 2: Comparing the response length with frontier LRMs.
To further evaluate the reasoning efficiency of LRMs trained with REO-RL (Exp) in comparison to advanced LRMs, we conduct a controlled analysis focused on the correct response length. Since frontier LRMs and those fine-tuned with REO-RL (Exp) differ significantly in overall accuracy, we construct a balanced subset of 71 questions from the test set, where all models achieve an accuracy exceeding 50%. For each model, we compute the average length of correct reasoning traces for each problem. The final length metric for each model is then obtained by averaging the correct response lengths across all 71 questions. As shown in Tab. 2, the model obtained through fine-tuning DeepSeek-R1-Distill-Qwen-7B with REO-RL (Exp) exhibits more concise reasoning patterns than frontier LRMs, with a much shorter response length.
6.4 Ablation Study of REO-RL
Method AIME 2024 AIME 2025 AMC 2023 Minerva Math Average
Acc $\uparrow$ Length $\downarrow$ REG Acc $\uparrow$ Length $\downarrow$ REG Acc $\uparrow$ Length $\downarrow$ REG Acc $\uparrow$ Length $\downarrow$ REG Acc $\uparrow$ Length $\downarrow$ REG
Base LRM 55.3 13062.1 1887.4 39.7 14241.9 1527.0 90.9 6177.3 1611.4 43.1 5575.8 855.7 57.2 9764.3 1470.4
Vanilla RL 66.2 14264.7 1579.0 52.9 16305.3 1453.1 93.9 7259.8 1722.2 44.6 7300.7 945.2 64.4 11282.6 1424.9
REO-RL (Exp) 64.0↓2.3 7671.5 $\downarrow{60.7\%}$ 48.8↓4.2 8361.1 $\downarrow{57.3\%}$ 93.4↓0.5 4144.6 $\downarrow{48.4\%}$ 45.5↑0.9 3687.0 $\downarrow{59.4\%}$ 62.9↓1.5 5966.0 $\downarrow{55.9\%}$
REO-RL (Oracle) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 $4$.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
REO-RL (Task-Specific) 63.9↓2.4 8407.8 $\downarrow{62.0\%}$ 51.4↓1.6 9298.5 $\downarrow{75.3\%}$ 93.4↓0.5 4230.6 $\downarrow{54.5\%}$ 43.9↓0.7 3581.0 $\downarrow{38.9\%}$ 63.1↓1.3 6379.4 $\downarrow{59.3\%}$
REO-RL (Task-Specific-Hard) 61.5↓4.8 6804.2 $\downarrow{53.5\%}$ 47.9↓5.0 7092.5 $\downarrow{79.8\%}$ 92.0↓2.0 3253.7 $\downarrow{60.7\%}$ 44.9↑0.3 2314.7 $\downarrow{72.8\%}$ 61.5↓2.9 4866.3 $\downarrow{65.6\%}$
REO-RL (Linear) 63.5↓2.7 8982.0 $\downarrow{46.1\%}$ 49.9↓3.0 9001.6 $\downarrow{58.8\%}$ 93.90.0 4290.9 $\downarrow{48.4\%}$ 44.9↑0.3 3733.6 $\downarrow{47.5\%}$ 63.1↓1.3 6502.0 $\downarrow{50.3\%}$
REO-RL (Exp) - K=10 64.0↓2.3 9352.7 $\downarrow{45.2\%}$ 48.9↓4.1 9804.9 $\downarrow{47.8\%}$ 94.1↑0.2 4779.6 $\downarrow{52.0\%}$ 44.9↑0.3 4151.8 $\downarrow{46.3\%}$ 63.0↓1.5 7022.3 $\downarrow{48.1\%}$
REO-RL (Exp) - Coef=1 62.6↓3.6 7218.5 $\downarrow{70.0\%}$ 47.2↓5.7 7525.9 $\downarrow{73.0\%}$ 93.1↓0.8 3736.4 $\downarrow{63.4\%}$ 45.4↑0.8 3052.7 $\downarrow{73.4\%}$ 62.1↓2.3 5383.4 $\downarrow{69.3\%}$
Table 3: Results of Ablation Study on DeepSeek-R1-Distilled-Qwen-7B on All Benchmarks
We conduct an ablation study on the design choices of REO-RL using the DeepSeek-R1-Distill-Qwen-7B model to better understand its flexibility and performance characteristics. We demonstrate that REO-RL can maintain competitive performance across different configurations.
Token budget selection strategy. We evaluate REO-RL (Oracle) that adopts the oracle greedy strategy. This variant achieves slightly better performance than REO-RL (Exp), reflected in a lower REG. We also investigate linearly spaced token budgets, which leads to less optimal efficiency and higher REG.
Coefficient $c_{i}$ in REO-RL objective. Setting all coefficients $c_{i}$ uniformly to 1 leads the model to align more closely with the efficiency frontier and shorter response lengths. However, this approach comes at the cost of reduced overall accuracy.
Number of selected token budgets $N$. We increase the number of selected token budgets to $N=10$ to explore whether a finer granularity improves performance. In practice, we observe that using more token budgets results in slower convergence and a higher rate of training instability. Consequently, this configuration produces weaker results overall.
Question-Specific Oracle Budgets. We investigate a question-specific oracle strategy that assigns two token budgets to each problem, the minimum reasoning token budget derived through the frontier estimation experiments (Sec. 4.2) and the full budget $L_{\text{max}}$. This oracle approach performs better than REO-RL (Exp) on DeepSeek-R1-Distill-Qwen models but leads to less REG reduction than REO-RL (Exp) on Qwen3 models.
7 Conclusion
In this work, we investigate efficient reasoning for LRMs. We introduce reasoning efficiency frontiers, which characterize the empirically optimal trade-off between response length and accuracy for LRMs. To quantify the reasoning efficiency of a fine-tuned LRM, we introduce the Reasoning Efficiency Gap (REG), a unified metric that captures both accuracy and length. We benchmark existing methods and reveal a substantial gap between current fine-tuning approaches and the frontiers. Our proposed method, REO-RL, could largely enhance the reasoning efficiency and become the closest to the optimal length-accuracy trade-off.
Appendix A Additional Discussion
Connection to MRT [33] and BRPO [33].
Similar to our approach, REO-RL, both MRT and BRPO share the idea of optimizing rewards under limited token budgets. However, key differences exist between REO-RL and these prior or concurrent approaches.
[…]
Second, the ways to select partial responses are different in all of these works. MRT relies on partitioning the model-generated responses into semantically complete steps through keywords. BRPO selects a small set of linearly spaced token budgets. REO-RL introduces varying types of token budget selection strategies, such as exponentially spaced token budgets that achieve better results than the linear counterpart, and applying question-specific token budgets. Finally, MRT optimizes the single-step rewards while both BRPO and REO-RL employ dense reward training and optimize total rewards across a wide range of token budgets.
Appendix B Reproducibility
We will provide our code in the https://github.com/samjia2000/Optimal-Reasoning-Efficiency. Please refer to Sec. F for the details on reasoning efficiency frontiers and reasoning efficiency gap, and Sec. C for the implementation details.
Appendix C Implementation Details
Training Data.
For training data, we integrate data from DeepScaleR [26] and AReaL [35]. For models with at least 4B sizes, we use the training data of AReaL-Boba-RL-7B [35]. For 1.5B, we adopt the mixture of training data from DeepScaleR [26] and AReaL [35] and remove duplicated problems.
We implement the training algorithm with the AReaL framework [35], which supports SGLang [60] for rollout generation. Below we detail implementation details of each training algorithm.
Online RL Training.
We use PPO as the default online RL algorithm. Following standard practices in RL training for LLM reasoning [55, 13], we do not utilize value model and KL regularization. The default training setting and hyperparameters for PPO training are listed in Tab. 4.
For REO-RL and baseline methods, we do not carry out RL training directly from the base LRM since it would lead to prolonged training and slower convergence. Instead, we perform further fine-tuning on the RL trained versions for both 1.5B and 7B settings. Specifically, we adopt AReaL-Boba-RL-1.5B [35] and Skywork-OR1-Math-7B [11] as the starting points for further fine-tuning. Following the cluster configuration in Tab. 4, each experiment could finish within 4K GPU hours.
Table 4: Default training configurations and hyperparameters for PPO.
Training Configuration
Batch size (number of prompts) 128
Rollouts per prompt 16
Random seed 1
Cluster Config $8\times 8$ H800 (for 1.5B) / $16\times 8$ H800 (others)
PPO Parameters
PPO Minibatches 4
Clipping $\epsilon$ 0.2
Advantage normalization True
Discount factor $\gamma$ 1.0
GAE $\lambda$ 1.0
Epochs 2.0
Optimizer Parameters
Optimizer Adam
Learning rate $5\times 10^{-6}$
Weight decay 0.05
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Adam $\epsilon$ $1\times 10^{-5}$
Gradient norm clipping 1.0
Learning rate scheduler constant
Warmup steps proportion 0.001
Generation Parameters
Temperature 1.0
Top-p 1.0
Top-k -1
Max prompt length 1024
Min generation length 0
Max generation length 24376 (for 1.5B) / 32768 (others)
Supervised Fine-Tuning.
The default training configurations and hyperparameters for SFT are listed in Tab. 5.
Table 5: Default training configurations and hyperparameters for SFT.
Training Configuration
Batch size (number of prompt-answer pairs) 512
Cluster Config $16\times 8$ H800
SFT Parameters
Epochs 10
Save Frequency Steps 100
use_bf16 True
Max Seq Length 32768
Optimizer Parameters
Optimizer Adam
Learning rate $1\times 10^{-5}$
Weight decay 0.05
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Adam $\epsilon$ $1\times 10^{-5}$
Gradient norm clipping 1.0
Learning rate scheduler constant
Warmup steps proportion 0.03
Preference Learning.
We implement SimPO [28] in the AReaL framework [35]. The default training configurations and hyperparameters for SimPO are listed in Tab. 5.
Table 6: Default training configurations and hyperparameters for SimPO.
Training Configuration
Batch size (number of preference pairs) 128
Cluster Config $16\times 8$ H800
SimPO Parameters
Epochs 2
Save Frequency Steps 10
use_bf16 True
Max Seq Length 32768
SimPO Coefficient $\beta$ 1/2
SimPO Coefficient $\gamma$ 1.2/1.4
Optimizer Parameters
Optimizer Adam
Learning rate $1\times 10^{-5}$ (for 1.5B) / $3\times 10^{-6}$ (or 7B)
Weight decay 0.05
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Adam $\epsilon$ $1\times 10^{-5}$
Gradient norm clipping 1.0
Learning rate scheduler constant
Warmup steps proportion 0.03
Appendix D Baselines
RL with Token Budgets.
In our online RL training with token budget constraints, we control the maximum generation length during each training phase. Rather than fine-tuning the LRM directly on a fixed token budget, we adopt a progressive length-shrinking strategy. We begin training with a 16K token budget. Once RL training at this level converges, we reduce the budget to 8K and continue training. This process is repeated, halving the token budget each time, until we reach the minimum budget of 512 tokens. This staged approach enables the model to gradually adapt to shorter generation lengths while maintaining reasoning performance.
RL with Length Rewards.
In the “RL with Length Group Normalized Rewards” baseline [2], for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
$r(x,y^{i})$ $=\mathbb{I}\{y^{i}\text{ is correct}\}(1-\alpha f(|y^{i}|))$
where the function $f$ normalizes $|y^{i}|$ according to the lengths of correct responses and applies a sigmoid function. Specifically,
$f(|y^{i}|)=\sigma\left(\frac{|y^{i}|-\texttt{MEAN}(x)}{\texttt{STD}(x)}\right)$
where
$\texttt{MEAN}(x)$ $=\mathbb{E}_{y\sim\pi(\cdot|x),s.t.y\text{ is correct}}[|y|]$
$\texttt{STD}(x)$ $=\sqrt{\mathrm{Var}_{y\sim\pi(\cdot|x),s.t.y\text{ is correct}}[|y|]}$
In the “RL with Length-Harmonizing Rewards” baseline [24], for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
$r(x,y^{i})$ $=\frac{\overline{L}_{ref}(x)}{|y|}-1+\gamma\cdot(\mathbb{I}\{y\text{ is correct}\}-\overline{A}_{ref}(x))$
where $\overline{L}_{ref}(x)$ is the average response length of the reference model when taking $x$ as input and $\overline{A}_{ref}(x)$ is the average accuracy of the reference model. In our implementations, we set the models that serve as the starting point of RL training as the reference models.
In the MRT baseline [34], different from the original paper that only implements single-step optimization with offline collected response prefixes, we implement the online RL training version with dense rewards for MRT. For each question $x$ and the corresponding set of sampled responses , we partition each response into several steps $y=(y_{1},\cdots,y_{s})$. In each training step, the model is updated by computing the policy gradient for the following objective,
$\mathbb{E}_{x,y=(y_{1},\cdots,y_{s})\sim\pi_{\theta^{\prime}}(\cdot|x)}[$ $\sum_{i=1}^{s}\mathbb{E}_{y^{\prime}_{i}\sim\pi_{\theta}(\cdot|x,y_{:i-1})}[\mathcal{R}(x,\text{Answer}(\pi_{\theta^{\prime}},x,[y_{:i-1};y^{\prime}]))$
$-\mathcal{R}(x,\text{Answer}(\pi_{\theta^{\prime}},x,y_{:i-1}))+\alpha\cdot\mathcal{R}(x,y)]]$
where $\alpha$ is the weight for the overall accuracy and is set as $0.2$.
Hybrid Reasoning.
We use “\$\backslash n$” and “\$\backslash n$\$\backslash n$</think>” as the tokens to start thinking and no-thinking. We first run SFT over model-generated responses in the thinking and no-thinking modes to ensure the model initially has a non-zero probability for no-thinking modes. Then we apply HGPO training with a margin of $0.2$ following  [15].
Supervised Fine-Tuning.
In $\text{SFT}_{\text{Shortest}}$, we generate $16$ outputs for each question in the training dataset. Then we select the correct response with the shortest length for each question to construct the SFT dataset. In $\text{SFT}_{\text{TOPS}}$, we follow [51] to prompt the LRM to generate responses with three different types of reasoning efforts. We strictly follow the prompts used in [51]. For each type of reasoning effort, we generate $16$ responses. To construct the SFT dataset, the shortest correct response among all $48$ responses are gathered.
Preference Learning.
We adopt three strategies for constructing the preference datasets. In $\text{SimPO}_{\text{Shortest}}$, we adopt the responses generated for $\text{SFT}_{\text{Shortest}}$ and select the shortest correct response and the longest response as the preference pair for each question. In $\text{SimPO}_{\text{TOPS}}$, we use the same preference construction strategy as $\text{SimPO}_{\text{Shortest}}$ but on the responses generated through TOPS, which contain reasoning traces with different reasoning efforts. Finally, in $\text{SimPO}_{\text{DAST}}$, we again utilize the responses generated for $\text{SFT}_{\text{Shortest}}$ but adopt a different preference construction strategy. In the preference dataset of $\text{SimPO}_{\text{DAST}}$, each pair falls into one of the two cases: it either contains two correct responses where the positive sample is much shorter than the negative sample, or contains two incorrect responses where the positive sample is much longer than the negative sample.
Other Methods We Have Tried.
We have also experimented with Z1 [56] that constructs code-based reasoning traces, and L1 [1] that fine-tunes the LRMs to follow token budget instruction with RL. We find Z1 having poor performance on mathematical reasoning tasks, demonstrating significantly lower accuracy to that of the base LRMs as illustrated in Tab. 7. For L1, we find L1 mainly works under tight token budgets, i.e. less than $4K$. When we extend the training approach of L1 to a larger context length, i.e. 24K for 1.5B models, we find it hard to make the LRM learning to follow strict token budget instructions through RL. Consequently, the resulted models, L1-Exact-24K-1.5B and L1-Max-24K-1.5B could not follow the token budget instruction, as shown in Tab, 8.
Method AIME24 MATH500 GPQA
Accuracy (%) Length Accuracy (%) Length Accuracy (%) Length
Base LRM 31.5 16747.6 83.6 5633.1 44.6 10325.2
Z1 10.0 15106.0 63.6 4904.1 61.6 9004.1
Table 7: Result of Z1 on DeepSeek-R1-Distill-Qwen-1.5B. [56]
Instructed Token Budget AMC23 AIME24 AIME25
Accuracy (%) Length Accuracy (%) Length Accuracy (%) Length
2048 63.9 20267.7 38.6 18067.4 27.2 19657.7
4096 63.5 20142.6 38.4 17961.3 26.9 19711.1
8192 62.9 20191.6 38.4 17868.7 26.8 19673.5
Table 8: Result of L1-Exact-24K-1.5B.
Appendix E REO-RL
E.1 Implementing REO-RL
Generation Phase.
In the generation phase of REO-RL, there are two rounds of LRM generation. In the first round, multiple responses are generated for each question in the training batch. In the second round, to compute the length-constrained rewards for each of the responses and across all selected token budgets, we choose all truncated responses $y_{:L_{i}}$ and apply a prompt to enforce the LRM to generate the final answer given incomplete reasoning traces, i.e. $a=\pi_{\theta}(\cdot|x,y_{:L},[\text{The Final Answer is}]).$ We follow the prompt employed by [7] and [8].
Prompt for Forcing LRM to Produce Answer
Oh, I suddenly got the answer to the whole problem. **Final Answer**:
[…]
Dense Reward RL.
REO-RL obtains dense rewards through forcing the LRM to generate answer under various token budgets. The objective of REO-RL is as follows,
REO-RL: $\mathscr{L}_{{REO-RL}}(\theta,\mathcal{D})=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[\sum_{i=1}^{N+1}c_{i}\mathcal{R}(x,\text{Answer}(\pi_{\theta},x,y_{:L_{i}}))\right]\right]$
where $c_{i}=\frac{L_{i+1}-L_{i-1}}{2}$ for $1\leq i\leq N$ and $c_{N+1}=\frac{L_{\text{max}}-L_{N}}{2}$ are the coefficient for the $i$-th token budget.
To perform policy update, we compute the return for each section between two consecutive token budgets $L_{i}$ and $L_{i+1}$. For $1\leq i\leq N$, we compute,
$\mathrm{Return_{i}(x,y)}=\sum_{j=i}^{N+1}c_{j}\mathcal{R}(x,\text{Answer}(\pi_{\theta},x,y_{:L_{j}}))$
Since we disable value model in the training process, the computed returns are then used directly as the advantages for loss computation. In practice, for each prompt, we apply group normalization for each $i\in[1,N]$ before the policy update step, i.e. $\mathrm{Adv}_{i}(x,y^{j})=\frac{\mathrm{Return(x,y^{j})}-\frac{1}{M}\sum_{j^{\prime}}\mathrm{Return}(x,y^{j^{\prime}})}{\sqrt{\mathrm{Var}_{j^{\prime}}[\mathrm{Return}(x,y^{j^{\prime}})]}}$. Then $\text{Adv}_{i}(x,y)$ will be used as advantages to update the tokens $y_{L_{i}:L_{i+1}}$.
E.2 REO-RL (Q-Spec)
REO-RL (Q-Spec).
Alternatively, we note that there should exist a specific minimum token budget $L^{x}$ for each question $x$. This minimum token budget reflects the lowest token budget under which the base LRM could learn to achieve the same accuracy as the well-trained counterpart under the full token budget.
To derive the question-specific minimum token budgets, we take a practical estimation approach that extracts from the rollouts generated during the training process of various configurations in Sec. 4.2. For a question $x$, the RL experiments in Sec. 4.2 produce a large amount of rollouts generated by different fine-tuned models associated with the corresponding length-constrained rewards, i.e. $\mathcal{D}_{\text{rollouts}}=\{(\hat{\theta},x,y^{1},\cdots,y^{M},r(x,\cdot;\hat{\theta}))\}$ where $\hat{\theta}$ denotes a fine-tuned LRM and $y^{1},\cdots,y^{M}$ are $M$ rollouts generated by $\hat{\theta}$ for question $x$. From $\mathcal{D}_{\text{rollouts}}$, we compute $L^{x}$ that satisfies,
$L^{x}=\min\left\{L|L\in[1,L_{\text{max}}]\text{ s.t. }\max_{(\hat{\theta},x,y^{1},\cdots,y^{M})\in\mathcal{D}_{\text{rollout}}}\frac{\sum_{i}r(x,y^{i}_{:L};\hat{\theta})}{M}\geq\max_{(\hat{\theta},x,y^{1},\cdots,y^{M})\in\mathcal{D}_{\text{rollout}}}\frac{\sum_{i}r(x,y^{i}_{:L_{\text{max}}};\hat{\theta})}{M}\right\}$ (13)
In practice, evaluating $r(x,y^{i}_{:L};\hat{\theta})$ for all history models $\hat{\theta}$, history rollouts and all possible $L$ could be expensive due to additional inference runs to evaluated truncated responses as in Eq. 4.1. When computing the minimum token budgets, we directly evaluate the correctness of the partial response $y^{i}_{:L}$ without the additional operation to produce the answer.
In REO-RL (Question-Specific), we set $N=1$ and $L_{1}=L^{x}$ for each question $x$, leading to the objective,
REO-RL (Q-Spec): $\mathscr{L}_{{REO-RL}\text{ (Q-Spec)}}(\theta,\mathcal{D})=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[c_{1}\cdot r(x,y_{:L^{x}};\theta)+c_{2}\cdot r(x,y_{:L_{\text{max}};\theta})\right]\right]$ (14)
where $c_{1}=\frac{L^{x}}{2},c_{2}=\frac{L_{\text{max}}-L^{x}}{2}$ following Eq. 14.
Appendix F Reasoning Efficiency Frontiers & Reasoning Efficiency Gap
Reasoning Efficiency Frontiers.
The two boxes below record the detailed lengths and accuracies for points on the estimated reasoning efficiency frontiers for DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B, respectively.
Reasoning Efficiency Frontier for Qwen3-4B x, y = [0, 64, 128, 192, 256, 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960, 1024, 2048, 3072, 4096, 5120, 6144, 7168, 8192, 9216, 10240, 11264, 12288, 13312, 14336, 15360, 16384, 17408, 18432, 19456, 20480, 21504, 22528, 23552, 24576, 25600, 26624, 27648, 28672, 29696, 30720, 31744, 32768], [0.05864545036764705, 0.07844286151960785, 0.08253484987745098, 0.09157475490196078, 0.09366000306372549, 0.10081380208333333, 0.10913181678921569, 0.12006548713235293, 0.14239813112745098, 0.16250957414215686, 0.18545879289215683, 0.20630935968137254, 0.22248008578431372, 0.23359183517156862, 0.24719860600490196, 0.25932329963235295, 0.26839958639705885, 0.37697610294117645, 0.4535807291666667, 0.5027018229166667, 0.5349207261029412, 0.5595760569852941, 0.5795496323529412, 0.5964384191176471, 0.6082050398284314, 0.6200099571078431, 0.6309244791666666, 0.6428308823529412, 0.6487189797794117, 0.6562059589460785, 0.6632372089460784, 0.6686408547794117, 0.6734145220588235, 0.6795764399509804, 0.6835171568627451, 0.6894416360294119, 0.693035768995098, 0.6970358455882353, 0.7003484987745099, 0.7034734987745098, 0.7032781862745098, 0.7061714920343137, 0.7074371936274509, 0.7081533394607843, 0.7090647977941177, 0.7103381587009804, 0.7101064644607843, 0.707257199754902]
Reasoning Efficiency Frontier for Qwen3-8B x, y = [0, 64, 128, 192, 256, 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960, 1024, 2048, 3072, 4096, 5120, 6144, 7168, 8192, 9216, 10240, 11264, 12288, 13312, 14336, 15360, 16384, 17408, 18432, 19456, 20480, 21504, 22528, 23552, 24576, 25600, 26624, 27648, 28672, 29696, 30720, 31744, 32768], [0.09000842524509803, 0.10032935049019608, 0.09675245098039215, 0.11996208639705883, 0.1271771599264706, 0.1278818167892157, 0.14018267463235295, 0.16002795649509804, 0.1714862898284314, 0.2006414675245098, 0.21497970281862744, 0.24749348958333334, 0.2608475030637255, 0.2683632046568627, 0.2772250306372549, 0.28248697916666665, 0.29307789522058825, 0.39007352941176476, 0.4767539828431373, 0.5257927389705882, 0.5632065716911765, 0.5868183210784313, 0.6074180453431373, 0.6239028033088235, 0.6389916513480391, 0.6473824295343137, 0.6582471660539215, 0.666616881127451, 0.6744542738970589, 0.6808746936274509, 0.6869829963235294, 0.6924479166666666, 0.695980775122549, 0.7023954503676471, 0.7065142463235294, 0.7109911151960785, 0.7135799632352942, 0.7167777267156862, 0.7191789215686275, 0.7239296109068628, 0.726351868872549, 0.7280005361519608, 0.7298598345588236, 0.7314587162990196, 0.7330135569852941, 0.7345473345588236, 0.7354587928921568, 0.7288985906862746]
Reasoning Efficiency Frontier for DeepSeek-R1-Distill-Qwen-1.5B x, y = [0, 64, 128, 192, 256, 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960, 1024, 2048, 3072, 4096, 5120, 6144, 7168, 8192, 9216, 10240, 11264, 12288, 13312, 14336, 15360, 16384, 17408, 18432, 19456, 20480, 21504, 22528, 23552, 24576, 25600, 26624, 27648, 28672, 29696, 30720, 31744, 32768], [0.06101600796568627, 0.05281671262254902, 0.06075367647058823, 0.07656441482843138, 0.09422679227941178, 0.09907322303921567, 0.11950635723039214, 0.14104051776960785, 0.16222426470588236, 0.17982153799019607, 0.19957299325980393, 0.22226179534313725, 0.23991076899509806, 0.25617723651960783, 0.2744064031862745, 0.2836722579656863, 0.29287109375, 0.3738262101715686, 0.413882506127451, 0.44010225183823526, 0.4490253523284314, 0.4542604932598039, 0.4631778492647059, 0.4701803768382353, 0.4771963082107843, 0.48008961397058825, 0.48356885723039217, 0.48675130208333334, 0.4884727328431372, 0.48968098958333334, 0.49039713541666663, 0.4909466911764706, 0.4908088235294118, 0.4905771292892157, 0.4908662683823529, 0.49117647058823527, 0.4919289981617647, 0.4914445465686274, 0.4919002757352941, 0.4920955882352941, 0.4919002757352941, 0.4919289981617647, 0.4919289981617647, 0.4919577205882353, 0.4919002757352941, 0.4919289981617647, 0.4920955882352941, 0.49215303308823527]
Reasoning Efficiency Frontier for DeepSeek-R1-Distill-Qwen-7B x, y = [0, 64, 128, 192, 256, 320, 384, 448, 512, 576, 640, 704, 768, 832, 896, 960, 1024, 2048, 3072, 4096, 5120, 6144, 7168, 8192, 9216, 10240, 11264, 12288, 13312, 14336, 15360, 16384, 17408, 18432, 19456, 20480, 21504, 22528, 23552, 24576, 25600, 26624, 27648, 28672, 29696, 30720, 31744, 32768], [0.06838809742647059, 0.07453469669117647, 0.08459520526960784, 0.09630629595588236, 0.09808900122549019, 0.11145450367647058, 0.14210707720588237, 0.16989123774509804, 0.1799383425245098, 0.2068627450980392, 0.22867455575980392, 0.24759880514705881, 0.26799938725490197, 0.2935891544117647, 0.31299019607843137, 0.33319546568627456, 0.38615196078431374, 0.4691272212009804, 0.50302734375, 0.5375919117647059, 0.5522518382352941, 0.5640356924019608, 0.5759803921568628, 0.5929974724264706, 0.6051547181372549, 0.61650390625, 0.6240272671568627, 0.6303423713235294, 0.63447265625, 0.63720703125, 0.6392252604166667, 0.64111328125, 0.64345703125, 0.6455403645833333, 0.6470377604166666, 0.6482747395833333, 0.6487955729166667, 0.6512044270833334, 0.6522460937500001, 0.6525065104166667, 0.6532877604166667, 0.6522460937500001, 0.6527669270833334, 0.6535481770833333, 0.6535481770833333, 0.6535481770833333, 0.6532877604166667, 0.6532877604166667]
Evaluating Reasoning Efficiency Gap.
To practically evaluate REG, instead of strictly following Eq. 6, we obtain approximations of $\sum_{L=1}^{L_{\text{max}}}\hat{J}_{\text{optimal}}$ and $\sum_{L=1}^{L_{\text{max}}}J(\mathcal{D},\theta,L)$ through numerical integration on a set of token budgets respectively, in a similar way to Eq. 9. We select $\{L_{1},\cdots,L_{N}\}=\{64i|0\leq i<16\}\cup\{1024i|1\leq i\leq 16\}$. Note that we set $L_{\text{max}}=16K$ instead of $L_{\text{max}}=32K$ to focus on the efficiency gap under lower token budgets.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We believe our introduction and the abstract sections both clearly state the contributions made in this paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to Sec. 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: N/A
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide full guideline for reproducing the results of our experiment results in Appendix. C, Appendix. D and Appendix. E.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide our training code through anonymous link in Appendix. B. Our training data are gathered from open-sourced projects, which we have already properly cited. We also provide the estimated reasoning efficiency frontiers, that are necessary for computing the introduced metric REG , in Appendix. F.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please refer to Sec. 6.1 and Appendix. C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We did not include error bars for end-to-end experiments because this paper includes a large volume of expensive experiments. We present results within a single trial across different settings.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please refer to Appendix. C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: N/A
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper investigates training algorithms, which has limited social impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We ensure existing assets are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the code for implementing REO-RL and the exact numerical value of our estimated efficiency frontiers. We will also release necessary models in the future.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This work does not use LLMs as any important, original, or non-standard components
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.